OpenStudy (cherie_magee):

5x + y = 6 5x + 3y = -4 The y-coordinate of the solution to the system shown is _____. -5 -1 1/2 1

OpenStudy (radar):

Are you ready to work on this problem?

OpenStudy (radar):

The "no response" indicates you are not ready. bye

OpenStudy (cherie_magee):

i was trying to figure it out o my own for a sec but yes im ready @radar

OpenStudy (radar):

Good, I see from your previous post, that you are using the substitution method to solve these "system of equations" problems. Do you want to use the substitution method for solving this?

OpenStudy (radar):

For this particular problem, I would proceed in this manner. Get the term 5x which is in both equations isolated (on one side of the equal sign) 

OpenStudy (radar):

Please respons, This is not supposed to be a monologue !!

OpenStudy (radar):



OpenStudy (radar):

Come back when you are ready to work this problem.

OpenStudy (radar):

In case you do come back, now that we have 6 -y = -3y - 4 3y-y = -4-6 2y = -10 y = -5